OpenStudy (anonymous):

find f(a+h) if f(x)=x^2-3x

OpenStudy (anonymous):

\[ f(dog) = dog^2 - 3 dog\\ f(dog+ cat)= (dog+cat)^2 - 3(dog+cat)\\ \] Can you finish it now?

OpenStudy (anonymous):

\[ f(a+h) =(a+h)^2 -3 (a+h) \]

OpenStudy (anonymous):

Expand and you are done

OpenStudy (anonymous):

yes that's perfect thanks!!!

OpenStudy (anonymous):

YW